OpenStudy (anonymous):

Mrs. Mitchell put a total of 10,000 into two accounts. One account earns 6% simple annual interest. The other account earns 6.5% simple interest. After 1 year, the two accounts earned 632.50 interest. Find how much money was invested in each account.

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

well if she puts x into 6% then she has to put 10000 - x into 6.5% so using the simply interest formula you have \[632.5 = 0.06 \times x + 0.065 \times (10000 - x)\] after disrtibuting and collecting like terms you get \[632.5 = 650 - 0.005x\] I'll let you solve for x... the amount invested at 6% when you get the answer... its 10000 - x for the other investment hope it makes sense

OpenStudy (anonymous):

3500 and 6500

OpenStudy (campbell_st):

thats it

OpenStudy (anonymous):

Thank you!